An urn contains 3 red balls, 4 white balls, and 5 blue balls. A ball is drawn at a random and its color is noted. A second ball is then drawn, without replacing the first ball, and its color is noted. Find the following probabilities.
(a) The first ball is red
(b ) The second ball is white given the first ball is red
(c ) The first ball is red and the second ball is white

Answer by sudhanshu_kmr(1152)


a) 3/12 = 1/4
b) 5/11
c) 3/12 * 5/11 = 5/44